# Teaching Human Behavior Improves Content Understanding Abilities Of LLMs

### v1

@@ 3.1 Evaluation
3.2 Discussion
Tables 1, 2, and 3 contain the results for the visual question answering tasks, Tables 4, 5, 6 contain the results for video and image understanding tasks, Tables 7 contains the results for dense-captioning, and Table 8 contains the results for image and video memorability benchmarks. A common trend we observe across all the results is that Behavior-LLaVA performs better than the base model LLaMA-Vid and the finetuned model Ad-LLaVA on all tasks, especially in zero-shot settings. Interestingly, the performance gains remain even after fine-tuning on the task dataset (Tables 4, 5, 6).
The performance gains are smaller for the more low-level tasks of action and object recognition (Tables 1, 2, and 3), but much higher for the more high-level tasks of emotion understanding, sentiment analysis, persuasion strategy classification, and memorability simulation. This indicates that receiver behavior has richer signals for higher-level tasks. The gains are observed across both image and video benchmarks. We also observe that classification using a story generated by GPT-3.5 (following Bhattacharyya et al. [13]) results in better performance than only using the video (Tables 1, 2, 4).
Figures 4, 5, 7, and 3, 6, 8, show several randomly sampled qualitative examples for dense captions generated by Behavior-LLaVA over images and videos respectively. It can be noticed that despite not being explicitly trained for this task, the model performs quite well, picking up various artistic, cognitive, and object and material properties. From Table 7, while Behavior-LLaVA shows a decrease in correctness over LLaMA-Vid, it shows significant improvement in other aspects, including detail and quality. On these aspects, it even comes close to 2.5X larger models (LLaVA-1.6 (34B)).
Next, we compare the signals from behavioral data of perception and action. For this, we compare Behavior-LLaVA trained on the perception data of Salicon using two instruction formats with Behavior-LLaVA trained on BLIFT. Further, within BLIFT, we compare the learnings from the task of predicting likes/views with the task of predicting only comments. Table 9 contains the results. It can be noted from the table that training on just Salicon results in a performance decrease for the lower-level task of action recognition on MSRVTT-QA but improves performance on the higher-level task of Emotion recognition. However, the gains are smaller than those observed with training on BLIFT.
4 Conclusion
Appendix A GPT-4 Prompts
Listing 4: GPT-4V Prompt to calculate correctness of a image dense caption
You are a great critique for analyzing images and captions.
Listing 5: GPT-4V Prompt to calculate detail of a image dense caption
You are a great critique for analyzing images and captions.
Listing 6: GPT-4V Prompt to calculate quality of a image dense caption
You are a great critique for analyzing images and captions.

### v3

@@ 4 Conclusion
Sampling Ratio Epoch Likes/Views $R^{2}$ Comments Perplexity Performance Base-Model 0 -0.1 6.22 0 1:1 0.5 0.11 4.71 5.49 1 0.22 3.95 8.23 1.25 0.33 3.19 10.97 1.5 0.35 3.13 11.79 2 0.38 3.08 12.31 2.2 0.4 3.05 12.57 1:2 0.5 0.14 4.33 3.04 1.05 0.28 3.66 7.08 1.45 0.42 2.99 8.12 1:10 0.5 0.15 3.43 1.38 0.8 0.31 2.78 3.44 1 0.38 2.46 4.48 1.2 0.46 2.13 5.51 2:1 0.5 0.1 5.3 3.52 1 0.21 4.6 7.45 1.5 0.31 3.9 9.13
Table 12: Ablation on different sampling ratios and epochs of training. Sampling ratio is the ratio of behaviour data to multimodal instruct data. Performance is the average increase in 0-shot accuracy on 6 tasks with 250 samples each from the eval set. These tasks include image emotion recognition, video emotion recognition, persuasion strategy classification, MSRVTT, HVU and MSVD-QA
Model Audio Summarization (3 Shot) IMDb Sentiment BLEU ROUGE METEOR 0-shot 1-shot Behaviour-LLaVA 19.0 25.1 39.3 84.1 90.2 LLaMA-VID 15.1 18.3 30.7 80.3 87.9 VALOR (Chen et al., 2023) 6.6 10.0 23.9 - - Improvement of Behavior-LLaVA over LLaMA-Vid 25% 37% 28.01% 4.73% 2.61%
Table 13: Evaluation on audio and text modalities. We evaluate on the audio summarization benchmark(Han et al., 2023) for audio and IMDB sentiment benchmark for text.(Maas et al., 2011)
Figure 11: Percentage performance improvement over an untrained LLaMA-Vid model, compared across various sampling ratios at different checkpoints. The 1:1 sampling ratio shows the best empirical performance. Performance is averaged over 0-shot accuracy improvements on six tasks with 250 samples each from the evaluation set. These tasks include image emotion recognition, video emotion recognition, and persuasion strategy classification, MSRVTT, HVU and MSVD-QA. The figure also shows the benefit of our data filtering process. While training on unfiltered BLIFT also improves the result over baseline performance of Llama-Vid but data filtering adds more improvement on top of it.
Appendix A Dataset Descriptions
  1. 1.
    MSVD-QA and MSRVTT-QA: These datasets are based on Microsoft Research Video Description (Chen & Dolan, 2011) and MSR-VTT corpora (Xu et al., 2016b) and are extensively used in many video captioning and question-answering experiments. The MSVD-QA dataset has a total number of 1,970 video clips and 50,505 question-answer pairs. The MSRVTT-QA dataset contains 10K video clips and 243k question-answer pairs.
  2. 2.
    ActivityNet-QA (Caba Heilbron et al., 2015) is a benchmark primarily for human activity understanding containing 849 hours of video, including 28,000 action instances.
  3. 3.
    VideoEmotion-8 (Asur & Huberman, 2010) dataset comprises 1,101 user-generated videos sourced from YouTube and Flickr, each containing a minimum of 100 videos per emotional category, as per Plutchik Wheel’s emotion model.
  4. 4.
    Ekman-6 (Xu et al., 2016a) dataset is compiled from social websites, with each of its 1,637 videos labeled with a single emotion category based on Ekman’s psychological research.
  5. 5.
    CAER (Lee et al., 2019) dataset, sourced from TV shows, consists of 13,201 clips with an average sequence length of 90, each manually labeled with six basic emotions, aligning with the Ekman-6 dataset.
  6. 6.
    IAPSa (Mikels et al., 2005) is a subset of IAPS, following the Mikels model with eight emotion categories such as amusement, awe, contentment, excitement, anger, disgust, fear, and sadness. It consists of 395 affective images, marking the first visual emotion dataset with discrete categories.
  7. 7.
    Emotion6 (Peng et al., 2015) features 1,980 images sourced from Flickr, each labeled by 15 annotators according to the Ekman model, covering six emotion categories: happiness, anger, disgust, fear, sadness, and surprise.
  8. 8.
    EmoSet (Yang et al., 2023) encompasses a total of 3.3 million images, including 118,102 from social networks and artistic sources, evenly distributed across various emotion categories. Based on the Mikels model, EmoSet is categorized into eight emotion categories.
  9. 9.
    Abstract (Machajdik & Hanbury, 2010) exclusively consists of color and texture combinations without recognizable objects. Differing from the IAPS dataset where emotions often stem from identifiable objects, the abstract paintings dataset was peer-rated via a web survey, with each image rated approximately 14 times. It comprises 228 images spanning eight categories similar to those in IAPS.
  10. 10.
    Memento10k(Newman et al., 2020) is a short-term video memorability dataset comprising 10,000 video clips, with 900,000 human memory annotations recorded at various delay intervals. The video clips were, on average, 3s long.
  11. 11.
    VideoMem (Cohendet et al., 2019) is comprised of 10,000 soundless videos, each lasting 7 seconds, accompanied by memorability scores. Memorability was measured twice: first, shortly after viewing and again 24-72 hours later to capture both short-term and long-term memorability effects.
  12. 12.
    LaMem (Khosla et al., 2015) dataset is a short-term image memorability dataset comprising of 60000 images. The dataset contains scene-centric images, object-centric images and other types such as images of art, images evoking certain emotions, and other user-generated images.
  13. 13.
    SUN (Isola et al., 2011) dataset is a short-term image memorability dataset comprising $2$ images that were sourced from the SUN database.
  14. 14.
    MemCat (Goetschalckx & Wagemans, 2019) dataset is a short-term image memorability dataset comprising 10000 images. It consists of five broader memorability-relevant semantic categories (animal, sports, food, landscapes, vehicles), with 2K exemplars each, further divided into different subcategories (e.g., bear, pigeon, cat, etc. for animal). The images were sourced from existing image sets: ImageNet, COCO, Open Images Dataset, and SUN.
  15. 15.
    MediaEval (Kiziltepe et al., 2021) utilized publicly available links to short-form video clips, each averaging 6 seconds in duration, with both short-term and long-term memorability scores. Short-term memorability evaluations were conducted on videos viewed within the preceding few minutes, while long-term memorability assessments were based on videos viewed within the previous 24 to 72 hours.
  16. 16.
    LAMBDA (I et al., 2024) is a long-term memorability consisting of 2205 video ads collected over 1749 participants covering 276 brands. The average video length is 33 seconds and the videos are highly complex, consisting of audio, logos, fast-moving scenes, emotions, etc.
Appendix B Limitations
In this paper, we try to show the hypothesis that training on the behavior modality improves learning of the content modality. We train the models on comments and likes to show this. We test our models on multiple benchmarks and obtain positive results. While we try to cover a wide variety of tasks and while results do conclusively show that the hypothesis is true, yet, we can test on more benchmarks covering even more tasks. Similarly, we can show it with multiple models, other than LLaMA-Vid.
Appendix C Broader Impacts
Our paper talks about how behavioral training can positively impact content understanding of VLMs. We think this will be useful in various content understanding applications such as question answering, captioning, etc.
Appendix D Ethical Implications
This paper demonstrates that training on behavioral modalities enhances the learning of content modalities. Models trained on user interactions such as comments and likes were tested on multiple benchmarks and yielded positive results. While these findings present exciting opportunities for advancing content understanding in AI systems, they also raise important ethical considerations that must be carefully addressed.
1. No personally identifiable information (PII) is used to improve content understanding. Instead, aggregated behavioral data, including replays, likes, and comments, is utilized, ensuring user privacy. We have implemented rigorous anonymization and aggregation techniques to protect individual user identities.
2. We explicitly acknowledge the inherent biases that may exist in data sourced from social media platforms such as Reddit and YouTube. Despite our rigorous data filtering and cleaning processes, we recognize that the dataset may still be subject to demographic skews, self-selection bias, and algorithmic influences. These biases could potentially lead to uneven model performance across different user groups or reinforce existing societal biases. To mitigate these issues, we emphasize the importance of considering these broader implications when applying our model and interpreting its results.
3. We acknowledge the need for greater cultural diversity in our dataset. To address this, we plan to release our artifacts as open-source and encourage community contributions to incorporate multilingual and multicultural data. This could involve expanding the range of subreddits and YouTube channels included in our dataset, with the aim of capturing a more diverse and representative spectrum of receiver behavior across different cultural contexts. By taking this approach, we hope to enhance the ethical considerations and societal impact of our work, providing a more holistic view of behavioral patterns in various cultural settings. However, we also recognize that this approach may introduce new challenges, such as ensuring the quality and reliability of community-contributed data.